OpenStudy (anonymous):

A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.00s for the ball to reach its maximum height. Find (a) the ball initial velocity & (b) the height it reaches?

OpenStudy (anonymous):

h = Vi *t -1/2*g*t^2 Vf=0=Vi -g*t

OpenStudy (anonymous):

plug in t=3 s and you have two equations with two unknowns...

OpenStudy (anonymous):

if you use g=9.8 m/s^2 Vi will be in m/s, h will be in meters.

OpenStudy (anonymous):

would it 0.4591

OpenStudy (anonymous):

no. try again. use the second equation to find Vi. plug that into the first equation to find h.

OpenStudy (anonymous):

0.306

OpenStudy (anonymous):

how are you getting these numbers?

OpenStudy (anonymous):

v=u-gt ?

OpenStudy (anonymous):

final velocity = 0

OpenStudy (anonymous):

1.5 intial velocity?

OpenStudy (anonymous):

vi*3-.5*g*3^2? =1.5?

OpenStudy (anonymous):

find Vi first.